Grade 12th passPhysical Chemistry

If Equivalent weight of S in a SO2 is it that Equivalent weight of S in a so3 is-(1) 8×2÷3 (2) 8×3÷2(3) 8×2×3 (4) 2×3÷8

Aman
8 Years agoGrade 12th pass

2 Answers

Vikas TU
8 Years ago
Dear Student,
Equivalent weight= Molecular weight / valency
        =32/6 =16/3.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Trikut
6 Years ago
Answer is 8x2/3.
Explanation:
   Eq wt=8
   V factor=2
   So,  s=8x2/v factor of O I.e 3
   Hence answer is 8x2/3   
 Thanks
Cheers